FULL-DIMENSIONALITY QUANTUM CALCULATIONS OF ACETYLENE/VINYLIDENE ISOMERIZA-TION

\underline{ALEX BROWN}, J. M. BOWMAN, S. ZOU, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, GA 30322.

The isomerization of acetylene to vinylidene is examined theoretically in full-dimensionality (six degrees-of-freedom) using a new ab initio potential energy surface ^{a}. Eigenfunctions and eigenvalues of the exact Hamiltonian, for zero total angular momentum, are obtained using a series of novel truncation/recoupling procedures that permits calculations up to very high energies. The Hamiltonian is given in diatom-diatom Jacobi coordinates, with the choice H_{2}-C_{2} for the two diatoms to exploit the full permutational symmetry of the problem. By examining expectation values of the eigenfunctions, a number of states are clearly identified with vinylidene-like characteristics. Corresponding calculations are also done for C_{2}D_{2}. Full dimensional simulations of the photodetachment spectra of [C_{2}H_{2}]^{−} and [C_{2}D_{2}]^{−} are done (within the Franck-Condon approximation) and compared to the experimental ones.

aS. Zou and J. M. Bowman, Chem. Phys. Lett. 368, 421 (2003).